A grocer has a contract with a produce wholesaler that specifies that the wholesaler will supply the grocer with grapefruit t(d) If the lower specification of .75 lb. is not being met, estimate the fraction of grapefruits that weigh less than .75 lb.
Answer #1

Answer:

a)

Given,

Range = xbar +/- 3*(Rbar/d2)

substitute values

= 0.8267 +/- 3*0.11/1.693

= 0.8267 +/- 0.1949

= (0.6318 , 1.0216)

b)

Maximum weight = 1.0216

c)

No

Minimum weight = 0.6318

d)

Consider,

Rbar/d2 = 0.11/1.693 = 0.06497

Consider,

P(X < 0.75) = P(z < (0.75 - 0.8267)/0.06497)

= P(z < - 1.18)

= 0.1190001 [since from z table]

= 0.1190

= 11.90%